OpenStudy (anonymous):

2y equals 5 what does y equal ? @One098

OpenStudy (anonymous):

Try divide both sides by 2

OpenStudy (one098):

If you divide both sides by 2, you'll get your answer.

OpenStudy (anonymous):

\(\dfrac{2y}{2} = \dfrac{5}{2}\Longrightarrow y = \dfrac{5}{2}=\cdots~?\)

OpenStudy (anonymous):

so the answer will be a decimal

OpenStudy (one098):

Yes.

OpenStudy (anonymous):

@One098 is the answer 2.5

OpenStudy (one098):

Yes.

OpenStudy (anonymous):

thanks